OpenStudy (anonymous):

simplify 10/n - 5/n - 7/10

OpenStudy (anonymous):

what is \(10-5\) ?

OpenStudy (anonymous):

really, serious question, because that is all you need

OpenStudy (anonymous):

\[\frac{10}{n}-\frac{5}{n}=\frac{10-5}{n}=\frac{5}{n}\]

OpenStudy (anonymous):

so you can either leave it as \[\frac{5}{n}-\frac{7}{10}\] or else you can subtract to put it as one fraction